RAREST TREES IN PANAMA ACCORDING TO DIGITAL HERBARIUM

RECORDS

R. CONDIT, R. PÉREZ, S. LAO, S. AGUILAR

Along with the Missouri Botanical Garden (Tropicos), the New York Botanical Garden, the Global Biodiversity Information Facility, and the University of Arizona, CTFS has assembled known herbarium records from Panama that have been digitized. Among these are 135719 specimens of tree species for which latitude, longitude, province, and Latin name are recorded, including 2278 species.

Many of those species are rare. For example, 476 are known in only one province, with 151 of those in Chiriqui. There are 41 species known from just a single specimen anywhere in the country (Table 1).

Date: June 18, 2012.

Table 1: Tree species known from just a single herbarium record in Panama, along with the province in which they are known and the number of CTFS inventories in which they are known.

Latin Province CTFS Sites
Acalypha cuneata  Darien  0
Ardisia furfuracella  Chiriqui  0
Ardisia scheryi  Chiriqui  0
Ardisia solanacea  Colon  1
Ardisia subsessilifolia  Chiriqui  0
Ardisia whitei  Chiriqui  0
Caryodaphnopsis tomentosa Bocas del Toro  1
Ceiba rosea  Darien  0
Clethra coloradensis  Chiriqui  0
Coussarea austin-smithii  Bocas del Toro  0
Diplotropis purpurea  Colon  1
Dussia tessmannii  Panama  2
Elaeagia myriantha  Darien  0
Euterpe oleracea  Darien  0
Gustavia verticillata  Darien  0
Inga capitata  Colon  3
Meliosma grandifolia  San Blas  0
Meriania phlomoides  0
Miconia ampla  Bocas del Toro  0
Miconia brevitheca  Darien  0
Miconia colliculosa  San Blas  0
Miconia oraria  San Blas  0
Myrsine cubana  Bocas del Toro  0
Neosprucea grandiflora  Colon  0
Opuntia elatior  Herrera  1
Oreopanax compactus  Chiriqui  0
Oxandra venezuelana  Chiriqui  1
Peltogyne purpurea  Darien  0
Pera oppositifolia  Chiriqui  1
Phyllanthus skutchii  Colon  0
Pithecellobium lanceolatum Herrera  1
Psychotria oresbia  Darien  0
Quararibea pumila  Bocas del Toro  1
Raputia subsigmoidea  Colon  0
Schefflera latiligulata  Cocle  0
Schefflera seibertii  Bocas del Toro  0
Sideroxylon celastrinum  Herrera  1
Strychnos chlorantha  Darien  0
Telanthophora arborescens Chiriqui  0
Tetrorchidium rotundatum Veraguas  0
Zanthoxylum fagara  Panama  0

Figure 0.1. Locations of those 41 tree species known from single herbarium records in Panama